a bank loaned out $18,000, part of it at the rate of 6% and the rest at 14%. If the interest received in one year totaled $2,000, how much was loaned at 6%?
Answer by nerdybill(7384)

a bank loaned out $18,000, part of it at the rate of 6% and the rest at 14%. If the interest received in one year totaled $2,000, how much was loaned at 6%?
.
Let x = amount ($) loaned at 6%
then
18000-x = amount ($) loaned at 14%
.
.06x + .14(18000-x) = 2000
.06x + 2520-.14x = 2000
2520-.08x = 2000
-.08x = -520
x = $6500 (answer)